A certain type of ochro seed germinates 75% of the time. A backyard farmer planted 6 seeds. What is the probability that more than 4 seeds germinate?

1
Expert's answer
2022-02-06T15:16:14-0500

Let "X=" the number of seeds which germinate: "X\sim Bin (n, p)."

Given "n=6, p=0.75, q=1-p=1-0.75=0.25"

"P(X>4)=P(X=5)+P(X=6)"

"=\dbinom{6}{5}(0.75)^5(0.25)^{6-5}+\dbinom{6}{6}(0.75)^6(0.25)^{6-6}"

"=0.35595703125+0.177978515625"

"\approx 0.5339"